# Programming Every Example: Lifting Pre-training Data Quality Like Experts at Scale

### v1

@@ Abstract
Large language model pre-training has traditionally relied on human experts to craft heuristics for improving the corpora quality, resulting in numerous rules developed to date. However, these rules lack the flexibility to address the unique characteristics of individual example effectively. Meanwhile, applying tailored rules to every example is impractical for human experts. In this paper, we demonstrate that even small language models, with as few as 0.3B parameters, can exhibit substantial data refining capabilities comparable to those of human experts. We introduce Programming Every Example (ProX), a novel framework that treats data refinement as a programming task, enabling models to refine corpora by generating and executing fine-grained operations, such as string normalization, for each individual example at scale. Experimental results show that models pre-trained on ProX-curated data outperform either original data or data filtered by other selection methods by more than $2\%$ across various downstream benchmarks. Its effectiveness spans various model sizes and pre-training corpora, including C4, RedPajama-V2, and FineWeb. Furthermore, ProX exhibits significant potential in domain-specific continual pre-training: without domain specific design, models trained on OpenWebMath refined by ProX outperform human-crafted rule-based methods, improving average accuracy by $\mathbf{7.6\%}$ over Mistral-7B, with $\mathbf{14.6\%}$ for Llama-2-7B and $\mathbf{20.3\%}$ for CodeLlama-7B, all within $\mathbf{10}$B tokens to be comparable to models like Llemma-7B trained on $\mathbf{200}$B tokens. Further analysis highlights that ProX significantly saves training FLOPs, offering a promising path for efficient LLM pre-training. We are open-sourcing ProX with $\mathbf{\geq 100}$B corpus, models, and sharing all training and implementation details for reproducible research and future innovation.
@@ 1 Introduction
Experimental results demonstrate that the proposed ProX framework consistently lifts data quality for pre-training. Specifically, ProX achieves an average improvement of $2.1\%$ over $10$ downstream benchmarks and outperforms state-of-the-art data selection methods by over $2.0\%$. Furthermore, ProX shows broad applicability across model sizes from $0.3$B to $1.7$B and shows consistent performance gains across diverse pre-training corpora of varying quality, including RedPajama-V2 (Together, 2023), C4 (Raffel et al., 2020), and FineWeb (Penedo et al., 2024a). In domain-specific continual pre-training, ProX yields an $11\%$ gain over OpenWebMath (Paster et al., 2024) for TinyLlama-1.1B and $7.6\%$ for Mistral-7B across $9$ mathematical tasks, with similar improvements seen on Llama-2-7B and CodeLlama-7B. Beyond performance gains, results also suggest that pre-training on the refined corpus significantly boosts pre-training efficiency, achieving similar downstream performance with up to $\mathbf{20}\times$ less computing. We believe it is worthwhile to scale up computing FLOPs for data refinement, which enables similar performance with much less training cost and offers a promising path for efficient LLM pre-training.
@@ 3.1 Experiment Setup
We utilize various corpora for both general and specific domain data in our experiments. For general domain data, we begin with RedPajama-V2 (Together, 2023), a preprocessed large-scale dataset of $30$ trillion tokens from diverse Internet sources, ready for pre-training. We further apply ProX on the C4 corpus (Raffel et al., 2020) with $198$ billion tokens and the FineWeb dataset (Penedo et al., 2024a) containing $15$ trillion tokens, noted for high data quality. For specific domain experiments, we use OpenWebMath (Paster et al., 2024), a math-focused dataset with $15$ billion tokens. Given the limitations in computational resources, we conduct experiments on a randomly sampled subset of the entire pre-training dataset. See Table 7 (§ B.2) for sampling details.
@@
    In § 3.3, to test ProX on different model sizes and training corpora, we finally scale the TLM-m’s training tokens to $50$B over RedPajama-V2, C4, and FineWeb. To show ProX efficiency, we then directly compare with models covering a variety of pre-training approaches including (1) large-scale pre-training: TinyLlama-1.1B (Zhang et al., 2024b) trained on $3$T tokens; (2) model pruning from existing models: (SheadLlama (Xia et al., 2024) pruned from Llama-2 and trained on extra 50B tokens); (3) LLM synthesis (InstructionLM-1.3B (Cheng et al., 2024) trained on Mistral-7B generated data and cosmo-1.8B (Ben Allal et al., 2024) trained on Mistral-8x7B generated data).
@@ 3.3 Applying ProX across model sizes and pretraining corpora
Figure 6: Performance of original data and ProX curated data trained models across different datasets using $\approx 50$B tokens and comparison with existing models trained using different techniques. Inst-LM represents InstructionLM-1.3B; Cosmo represents Cosmo-1.8B; S-Llama represents ShearedLlama-1.3B.
@@
To assess the applicability of ProX across various pre-training corpora, we extend our experiments beyond RedPajama-V2 to include C4 (Raffel et al., 2020), and the recently released $15$-trillion-token pre-training corpus, FineWeb (Penedo et al., 2024a). For consistency, we apply exactly the same ProX-xs refining models detailed in Table 3.3 to these corpora without constructing new SFT data for each corpus. We conducted larger-scale experiments by training our model on approximately $50$ billion tokens, again achieving notable improvements. On ten downstream benchmarks, models trained on our method’s curated data showed improvements of $+2.0\%$ on RedPajama-V2, $+3.1\%$ on C4, and $+2.4\%$ on FineWeb.
@@ E.1 Case Studies
E.2 Computing Overhead Analysis

### v2

@@ Abstract
Large language model pre-training has traditionally relied on human experts to craft heuristics for improving the corpora quality, resulting in numerous rules developed to date. However, these rules lack the flexibility to address the unique characteristics of individual example effectively. Meanwhile, applying tailored rules to every example is impractical for human experts. In this paper, we demonstrate that even small language models, with as few as 0.3B parameters, can exhibit substantial data refining capabilities comparable to those of human experts. We introduce Programming Every Example (ProX), a novel framework that treats data refinement as a programming task, enabling models to refine corpora by generating and executing fine-grained operations, such as string normalization, for each individual example at scale. Experimental results show that models pre-trained on ProX-curated data outperform either original data or data filtered by other selection methods by more than $2\%$ across various downstream benchmarks. Its effectiveness spans various model sizes and pre-training corpora, including C4, RedPajama-V2, FineWeb, FineWeb-Edu, and DCLM. Furthermore, ProX exhibits significant potential in domain-specific continual pre-training: without domain specific design, models trained on OpenWebMath refined by ProX outperform human-crafted rule-based methods, improving average accuracy by $\mathbf{7.6\%}$ over Mistral-7B, with $\mathbf{14.6\%}$ for Llama-2-7B and $\mathbf{20.3\%}$ for CodeLlama-7B, all within $\mathbf{10}$B tokens to be comparable to models like Llemma-7B trained on $\mathbf{200}$B tokens. Further analysis highlights that ProX significantly saves training FLOPs, offering a promising path for efficient LLM pre-training. We are open-sourcing ProX with $\mathbf{\geq 500}$B corpus, models, and sharing all training and implementation details for reproducible research and future innovation.
@@ 1 Introduction
Experimental results demonstrate that the proposed ProX framework consistently lifts data quality for pre-training. Specifically, ProX achieves an average improvement of $2.1\%$ over $10$ downstream benchmarks and outperforms state-of-the-art data selection methods by over $2.0\%$. Furthermore, ProX shows broad applicability across model sizes from $0.3$B to $1.7$B and shows consistent performance gains across diverse pre-training corpora of varying quality, including RedPajama-V2 (Together, 2023), C4 (Raffel et al., 2020), FineWeb, FineWeb-Edu (Penedo et al., 2024a), and DCLM (Li et al., 2024). In domain-specific continual pre-training, ProX yields an $11\%$ gain over OpenWebMath (Paster et al., 2024) for TinyLlama-1.1B and $7.6\%$ for Mistral-7B across $9$ mathematical tasks, with similar improvements seen on Llama-2-7B and CodeLlama-7B. Beyond performance gains, results also suggest that pre-training on the refined corpus significantly boosts pre-training efficiency, achieving similar downstream performance with up to $\mathbf{20}\times$ less computing. We believe it is worthwhile to scale up computing FLOPs for data refinement, which enables similar performance with much less training cost and offers a promising path for efficient LLM pre-training.
@@ 3.1 Experiment Setup
We utilize various corpora for both general and specific domain data in our experiments. For general domain data, we begin with RedPajama-V2 (Together, 2023), a preprocessed large-scale dataset of $30$ trillion tokens from diverse Internet sources, ready for pre-training. We further apply ProX on the C4 corpus (Raffel et al., 2020) with $198$ billion tokens, the FineWeb dataset (Penedo et al., 2024a)(as well as FineWeb-Edu) containing $15$ trillion tokens, noted for high data quality, and DCLM-baseline-1.0 (Li et al., 2024). For specific domain experiments, we use OpenWebMath (Paster et al., 2024), a math-focused dataset with $15$ billion tokens. Given the limitations in computational resources, we conduct experiments on a randomly sampled subset of the entire pre-training dataset. See Table 7 (§ B.2) for sampling details.
@@
    In § 3.3, to test ProX on different model sizes and training corpora, we finally scale the TLM-m’s training tokens to $50$B over RedPajama-V2, C4, FineWeb (as well as FineWeb-Edu) and DCLM-baseline-1.0. To show ProX efficiency, we then directly compare with models covering a variety of pre-training approaches including (1) large-scale pre-training: TinyLlama-1.1B (Zhang et al., 2024b) trained on $3$T tokens; (2) model pruning from existing models: (SheadLlama (Xia et al., 2024) pruned from Llama-2 and trained on extra 50B tokens); (3) LLM synthesis (InstructionLM-1.3B (Cheng et al., 2024) trained on Mistral-7B generated data and cosmo-1.8B (Ben Allal et al., 2024) trained on Mistral-8x7B generated data).
@@ 3.3 Applying ProX across model sizes and pretraining corpora
Figure 6: Performance of original data and ProX curated data trained models across different datasets using $\approx 50$B tokens and comparison with existing models trained using different techniques like LLM data synthesis and direct model pruning.
@@
To assess the applicability of ProX across various pre-training corpora, we extend our experiments beyond RedPajama-V2 to include C4 and the recently released top-quality corpus including FineWeb, FineWeb-Edu, and DCLM. For consistency, we apply exactly the same ProX-xs refining models detailed in Table 3.3 to these corpora without constructing new SFT data for each corpus. We conducted larger-scale experiments by training our model on approximately $50$ billion tokens, again achieving notable improvements. On ten downstream benchmarks, models trained on our method’s curated data showed improvements of $+2.0\%$ on RedPajama-V2, $+3.1\%$ on C4, $+2.4\%$ on FineWeb, $\mathbf{+0.9\%}$ on FineWeb-Edu, and $\mathbf{+1.7\%}$ on DCLM.
@@ E.1 Case Studies
E.2 Error Analysis
As shown in Table 35, the failure ratio across both refining stages (document-level and chunk-level) and domains (General and Math) is remarkably low ($<0.5$%). This demonstrates that ProX’s refining tasks are well-suited for small models. Specifically, for the General domain, failure ratios are $0.04$% for document-level and $0.36$% for chunk-level refining, with an average of $3.7$ function calls per program in the chunk-level stage. For the Math domain, these ratios are $0.06$% and $0.11$%, respectively, with an average complexity of $2.7$ function calls at the chunk-level stage.
Despite the low failure rates, we observed two prevalent failure cases in ProX’s programs:
  1. 1.
    Repeated output or empty output: This occurs when a program inadvertently generates duplicate outputs or fails to produce any meaningful results. Such failures are typically linked to improper loop conditions or insufficient constraints in processing logic.
  2. 2.
    Non-existent target removal: In some cases, ProX’s programs attempt to remove a string or line that does not exist in the input data. This leads to incomplete execution or errors in the program output, particularly in datasets with irregular formats or unexpected variations.
Table 35: Failure ratio and average complexity (function calls) for ProX refining stages across domains.
Domain Failure Ratio (doc-level) Failure Ratio (chunk-level) Complexity (AVG. function calls)
General Domain 0.04% 0.36% 3.7
Math Domain 0.06% 0.11% 2.7
As shown in Table 36, we present two failure cases to illustrate instances of repeated output and non-existent target strings.
Table 36: Failure cases from RedPajama-V2 during applying ProX. “...” denotes omitted content due to limited space. The notation like [004] is used to indicate the line number.
Case 1: Repeated output (or Empty output)
... [004] P: 114 1. The problem statement, all variables and given/known data Mercury is poured into a U-tube as in Figure P15.18a….Basically I don’t understand why you would know to set the two volumes equal to each other? How do you know the volumes are the same? ... [007] Related Discussions Mechanical Engineering 6 Introductory Physics Homework 0 General Engineering 1 Introductory Physics Homework 2 Introductory Physics Homework 2
Output by ProX: remove_lines(start=1, end=1) remove_lines(start=6, end=6) remove_lines(start=7, end=7) remove_lines(start=7, end=7) remove_lines(start=7, end=7) remove_lines(start=7, end
Case 2: Non-existent target string
... [195] 18. Sathyamoorthi, C. R., Mbekomize, C., Mapharing, M., & Selinkie, P. (2018). The Impact of Corporate Governance on Working Capital Management Efficiency: Evidence from the Listed Companies in the Consumer Services Sector in Botswana. International Journal of Economics and Finance, 10, 135. https://doi.org/10.5539/ijef.v10n12p135 [196] 19. Vu, T. M. T., Tran, C. Q., Doan, D. T., & Le, T. N. (2020). Determinants of Capital Structure: The Case in Vietnam. Journal of Asian Finance, Economics, And Business, 7(9), 159-168. https://doi.org/10.13106/jafeb.2020.vol7.no9.159 ...
Output by ProX: # Analysis: this ‘source_str‘ can not be found in the original text  normalize(source_str="https://doi.org/10.13106/jafeb.2020.vol6.no2.53", target_str="")
E.3 Computing Overhead Analysis